OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (anonymous):

slop of the line will be 1 hence: y=x+m , 3=2+m and m=1 so the equation will be y=x+1